Sunday 02/05/12

9:40 A – 8:30P, 10:00P – 1:00A

  • Flip fly plates
  • try relabeling Dm0 antibodies with cy3B and 647 (first round labels still pretty concentrated protein.  Most of this is not antibody but we just reduce the dilution when adding to sample.
  • label Dm0 antibody with 488 and 750.
  • Confocal check slides
  • Josh fixed alignment of TIRF — was missing objective completely, yielding only scattered light from the objective.  750 laser still down.
  • Etch new slide of H3K4 abdA, ubx, ftz labeled (slide 44). 20 min rxn.  Good images by confocal.  Does not look to be etched at all on STORM scope.
  • Etching new slide S45 (13 intron) 26 minutes.  sample looks good, imaging on STORM2.
  • Pc sample looks great, 140 positions of cc14 and later gastrulated embryo programmed in. Top position too close to edge, sample loses oil and focus lock loses registration.  run canceled.
  • attempted to remove top positions from list and add oil back to slide.  with fresh oil focus lock regains connection, however slide has shifted and positions have been lost.
  • x-y stage controls frozen.  experiment aborted and system shutdown.
This entry was posted in Summaries, Uncategorized and tagged , . Bookmark the permalink.